The addition of 9.54 kJ of heat is required to raise the temperature of 225.0 g of a liquid hydrocarbon from 20.5oC to 45.0oC. What is the heat capacity of this hydrocarbon?


1
Expert's answer
2021-03-16T08:25:41-0400

"q = mcΔt \\

9540 = 225.0 \times s \times (45.0 -20.5) \\

c = \frac{9540}{5512.5} \\

c = 1.73 \;J/gC"